RHIC Running: A problem with the agsrhichome server has disabled outside access to the rhic web sites. This server has been powered down and ITD is looking into the problem.
Late Friday we moved to 93 bunches. This led to higher luminosity, but tune problems have kept us from realizing these gains.
A series of problems has made for rough running this past weekend. The noise source that causes blow-up in the Blue emittance returned on Friday night and then switched off again Saturday.
Tune changes from above the diagonal to below it have reverted to the earlier tune. Currently we do not have the Blue and Yellow tunes on opposite sides of the diagonal.
Scheduling issues:
- AGS Cold Snake cooldown began 31 December, a week ago.
- LINAC start-up of RF cavities 3-6 has begun, expected completion 16 January.
- Full Day Maitenance on Wednesday 16 January 2008 preparations for polarized protons.
- Au-31+ preparations in vacuum may be completed by 17 January 2008. We don't want to be competing with Polarized proton Set-up in the AGS at the time we are doing Au-31+ work.
- Set-Up times for 200GeV transverse vs 500GeV longitudinal polarized protons is 2 weeks vs 3 weeks.
- PHENIX has requested pp set-up begin 24 Jan. STAR would prefer that it begin 1 Feb.
- Yellow Q7 magnets need polarity changes that require 8-12 hours. Blue shunts also require work. ZDCs at 6 and 8 o'clock need to be moved. Don Bruno is ready for it. Steve Tepekian also needs to prepare a lattice. Hi-pot snakes and rotators followed by low current and high current tests.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |